# What is glacier sliding?

### v1

@@ 6.2 Threshold velocities for regularised-Coulomb sliding
In Weertman-type sliding two parameters ($C$ and $m$) control the relationship, but even in the simplest regularised-Coulomb relationship at least one further parameter is required ($u_{t}$ in our usage in Table 1) to set the transition point to bounded traction. Laboratory experiments over sediment indicate a transition to bounded traction occurs at around 50 m a-1 (Zoet and Iverson,, 2020) in a controlled setting, while the transition for cavitation over hard beds at small scales (<25 m) is more variable, and possibly as low as 5 m a-1 (Helanow and others,, 2021). Given that uniformly smooth soft beds of consistent material properties, or hard beds of consistent geometry, are somewhat unlikely, it becomes complicated to suggest a reasonable physically based $u_{t}$ value that can be universally applied and we therefore recommend avoiding arbitrary assumptions on its value. Joughin and others, (2019) set a threshold velocity of 300 m a-1 for Pine Island Glacier, Antarctica, without physical justification, which sets a clear spatial boundary for the region of bounded traction. As bounded traction significantly increases the sensitivity of outlet glaciers to climate forcing (Ritz and others,, 2015; Tsai and others,, 2015; Brondex and others,, 2019), further work determining the factors controlling the threshold velocity (Trevers and others,, 2024) is necessary, though we note that the increased complexity inherent to a regularised-Coulomb sliding may make it less versatile when applied universally (Section 6.5). We further note that in marine-terminating settings it may be difficult to distinguish whether the $u_{t}$ value used results in a good match with observations because of an intrinsic switch in processes beyond the threshold velocity, or due to the increasing influence of the ocean on subglacial hydrology, with ice velocity simply acting as a proxy for proximity to the coast.
6.3 Revisiting Iken’s bound
The assertion by Lliboutry, (1968), Iken, (1981), Schoof, (2005), Gagliardini and others, (2007) and others that sliding relationships of the form Eq. 1 must be bounded following Eq. 6 is based on two-dimensional topography comprised of compound sinusoids (a ‘washboard topography’, Fig. 10a). In these studies an upper size for subglacial cavitation is not theorised and so cavitation is implicitly assumed to hold at all scales. However, in realistic three-dimensional topography there are many more pathways for pressurised water to escape from cavities (Gimbert and others,, 2021; Helanow and others,, 2021) which may naturally taper the number of large cavities as they grade into relatively scarce subglacial lakes (Bowling and others,, 2019). If large topographic obstacles are present, but large-scale cavities are not concomitant, form drag may come to surpass ‘classical’ cavitation as the dominant sliding process in a given grid cell area (Fig. 10). At present, no studies outline the size-distribution of cavities that may occur with realistic three-dimensional bed geometries, but cavitation studies generally do not insinuate cavities larger than 15 m (e.g. Vivian and Bocquet,, 1973; Walder and Hallet,, 1979; Helanow and others,, 2021). Establishing this distribution and its interplay with form drag and grid cell area is an urgent area for future research. We suggest that the unsuitability of Iken’s bound in regions of large topographic obstacles may be the main reason for the continued efficacy of a Weertman-type sliding relationship in ice-sheet settings (e.g. Maier and others,, 2021) despite the questionable underlying assumptions (Section 3.2), if form drag around obstacles is unmitigated by large cavities bounding the potential traction of the area.
Figure 10: A spatial bound to Iken’s bound? a Cavitation (blue) occurring where topography is sufficiently small adapted from Schoof, (2005). b Cavitation can occur at a sufficiently small scale (yellow highlight) but will not occur in larger wavelength depressions. z is height from baseline and x is distance along flow. The thick black line is the glacier bed. $R$ represents a possible region in consideration for the sliding relationship. This figure is drawn in two dimensions for comparison with previous studies, but the same principles apply in three dimensions with realistic topography.
6.4 Does a unified basal sliding relationship exist? What would it look like?
A defining objective of much research into sliding is to uncover a ‘universal’ sliding relationship which can provide an accurate description of sliding processes across all settings, scales, and time periods with the frequently implicit assumption that this is controlled by a single tunable coefficient (henceforth single coefficient) and/or calculated subglacial water pressure (e.g. Minchew and Joughin,, 2020). Lamentably, we suggest that a simple universal sliding relationship defined by a single coefficient may not be a reasonable prospect, whether based on physical processes or heuristic modelling. The complex nature of myriad sliding sub-processes and their scale and setting dependence (Section 3) present no intrinsic physical reason why they should neatly fall within a single-coefficient relationship and it is not possible to control the shape of the available range of sliding relationships (Sections 3, 5) through a single parameter (Fig. 11). It follows that we do not hold that similarities in the sliding relationships for small-scale cavitation (Helanow and others,, 2021) and ring-shear sediment sliding (Zoet and Iverson,, 2020) indicate that a unified sliding relationship has been obtained (Minchew and Joughin,, 2020). Instead, the studies reviewed here point instead to a compound relationship that bears closer resemblance to descriptions of ice strain comprising several deformation mechanisms (e.g. Eq. 3 of Goldsby and Kohlstedt,, 2001). As a purely illustrative example, $\tau_{b}$ could, at a given setting and scale, be well represented by
$\tau_{b}=f_{w}(u_{b})+f_{rC}(u_{b})$ (11)
where $f_{w}$ and $f_{rC}$ are Weertman-type and regularised-Coulomb sliding respectively, each controlled by distinct $m$ and $C$ values.
Figure 11: The issue of unification. As for Fig. 1 but where a is the maximum traction for a bounded relationship, b is the threshold velocity for a bounded relationship, and c is the slope of an unbounded relationship or a switch to a bounded relationship.
However, as standard inversion procedures are ill-suited to determining multiple parameters simultaneously (Morlighem and others,, 2010) and as a compound sliding relationship (Eq. 11) would complicate implementation, the choice between a compound friction law (e.g. Eq. 3) and single-term equation from Table 1 is a compromise between a more realistic representation of friction and less practical numerical implementation. The path to a ‘unified’, or in more realistic terms ‘optimum’, sliding relationship could therefore come by disregarding process-based models for traction and the development of an efficient, transparent, and easily applicable method for determining the best parameters at a more setting-specific scale. Nonetheless, while this approach could address several pressing issues in glacier and ice-sheet modelling, it is still important to grasp the physical underpinnings to produce tractable predictions of sea level rise and to asses how a change in physical conditions can influence sliding. If a process-based sliding relationship is sought, we suggest de-emphasizing a single process in favor of the interplay between multiple sub-processes. Additionally, an optimum sliding relationship should conform to the following:
  • $\bullet$ A clear breakdown of which sliding sub-processes are considered, and the conditions under which a given sub-process becomes dominant for overall traction.
  • $\bullet$ Explicit inclusion of form drag.
  • $\bullet$ An assessment of scale dependence and the spatial range over which the relationship can be considered valid.
  • $\bullet$ An indication of the position of the inner-outer flow boundary, and if this will vary based on the conditions of the bed. This is not straightforward and may be difficult to determine, but transparent inclusion is preferable to omission.
A very broad range of methodologies have been used to investigate sliding (Section 3), but it may be that an empirical numerical-modelling approach (e.g. Helanow and others,, 2021), building on the results of more granular studies of sub-processes, is best suited to addressing the necessarily large parameter space inherent to this question over a variety of scales.
6.5 Implications for glacier and ice-sheet modellers
6.5.1 Which basal sliding relationship to use?
@@
No single sliding relationship accounts for the full breadth of processes or observations that comprise glacier and ice-sheet sliding (Sections 3, 5), making it challenging to recommend one single-tunable-coefficient sliding relationship. Pending future investigations, we make the following suggestions based solely on heuristic studies (Section 5), whose empirically-validated nature makes them more suitable for application to glacier and ice-sheet models.
We suggest that for ice sheets Weertman-type sliding (or equivalently, pseudo-plastic sliding), with a setting-dependent $m$, presents the least-bad option. Given information from the limited number of studies available (Table 2) lower $m$ values ($\leq 4$) are then better suited for rougher settings inferred to be predominantly hard-bedded and higher $m$ values ($\geq 5$) are better suited for regions where topography is extensively smooth and the setting is inferred to be largely soft-bedded. Therefore, while $C$ is the main object of the inversion procedure, $m$ may be assigned based on additional constraints, such as topographic roughness. Where applied, high $m$ values approximating regularised-Coulomb-type sliding should feature careful justification for their use, particularly given the influence of bounded traction (and implicitly unbounded sliding) on ice-sheet stability (e.g. Parizek and others,, 2013; Ritz and others,, 2015; Lippert and others,, 2024). The particular mapping of parameters in the pseudo-plastic relationship to easily understandable outcomes in $u_{b},\tau_{b}$ space is also preferable to a link to quantities such as till friction angle, which implies a greater physical meaning in the sliding relationship than may actually be the case, though the outcome is the same. Except for where a clear need is identified, regularised-Coulomb sliding may be less optimum due to its dependence on at least three tuning parameters, and its inability to represent unbounded traction (unless unrealistically high $u_{t}$ values are used).
@@
If hydrology is to be incorporated, we recommend a simplified implementation of Budd sliding as used in Choi and others, (2022) and set as the default for ISSM (Larour and others,, 2012), which is functionally the same as Weertman-type sliding but with $N$ included as a multiplication factor. It is possible that this represents a more physically based alternative to regularised-Coulomb sliding at marine terminating glaciers (Section 6.2). A simplified hydrological model as used in Choi and others, (2022) based on elevation above sea level may then be reasonable over annual timescales (Section 4), though we stress that this is a pressing area for future research – as is exact way in which $N$ should be incorporated within Eq. LABEL:eq:1.
Separately, it seems reasonable to suggest that Weertman-type sliding with $3\leq m\leq 4$ is a a fairly appropriate choice for most if not all valley glaciers, and particularly for studies covering a very large number of valley glaciers in a simple manner (Dehecq and others,, 2018; Cook and others,, 2023), with no contradictory evidence that we are aware of.
@@ 6.5.2 A note on terminology
We hope that this paper provides a clear description of the inner flow sub-processes that a sliding relationship will account for under ideal circumstances, and may also ease some of the angst experienced by modelling practitioners (ourselves included) in labelling bed properties. When we invert for basal properties we are not solely inverting for ‘roughness’ (e.g. Berends and others,, 2023), ‘till friction angle’ (e.g. Sutter and others,, 2021), or ‘friction’ at the ice-bed interface (e.g. Sergienko and Hindmarsh,, 2013) – we are likely inverting for all of these processes, and several more, simultaneously. We use ‘ideal circumstances’ here because the traction field is often one of the few tunable parameters in an ice-sheet model and so will in practice also account for so-called ‘compensating errors’ arising from incorrectly prescribed englacial temperature, ice viscosity, subglacial topography, and surface and basal mass balance (Berends and others,, 2023). We propose ‘basal traction’ as a least-bad descriptor here as its more abstract nature does not feature the same physical-process misunderstanding that comes with terms such as basal friction which appears to imply slip at the ice-bed interface rather than sliding as defined here. We further suggest that ‘sliding relationship’ is more appropriate than ‘sliding law’. Physical laws may hold up well at atomic or astronomical scales, but the overlapping nature of phenomena in Earth Sciences applications makes a single ‘law’ unrealistic. Similarly, ‘Weertman-type’ and ‘regularised-Coulomb-type’, or ‘unbounded’ and ‘bounded’ are preferable to ‘Weertman’ or ‘regularised-Coulomb’ which emphasise a physical basis that may not exist. Last, while we use ‘glacier sliding’ throughout this paper to remain in keeping with its common usage, but make a distinction with ‘slip’ to refer strictly to ice-bed displacement, ‘glacier basal motion’ (e.g. Waller,, 2001; Hubbard,, 2002; Law and others,, 2023) is a more precise term.
@@ 7 Conclusions
Glacier sliding is not a straightforward topic. Here, we have done our best to integrate existing understanding into a relatively straightforward framework. This is centred upon an inner flow (Fig. 5) comprised of potentially numerous sliding sub-processes (Fig. 4), which in ideal circumstances are accurately captured by the sliding relationship used in numerical models and which allows for the co-existence of soft- and hard-bed sliding, stick-slip and continuous sliding, and form drag. However, while this framework provides a pathway for incorporating these processes into an ‘optimum’ process-based sliding relationship, the existing range of heuristic studies (Table 2) likely indicates that a simple ‘unified’ sliding relationship controlled by a single coefficient is not a realistic outcome. Nonetheless, Weertman-type sliding, or simplified Budd sliding where hydrology is included, with an $m$ determined based on setting, preferably with the use of sensitivity studies, is a reasonable working assumption based on available observational evidence.
Despite obvious obstacles there are clear priorities for future work. Principally, determining a simple universally deployable methodology for determining $m$ across a wide range of settings and formally reconciling process-based studies with heuristic studies to understand the physical reasoning for the efficacy of the sliding relationships outlined in Section 5. Overall, we hope that the framework presented here allows for a transparent approach to understanding glacier sliding, applying glacier sliding relationships to numerical models, and for directing future research.
Appendix A Appendix A
Figure 12: Appendix A. The difference between Zoet and Iverson, (2020) and Helanow and others, (2021) at high water pressures.. Zoet and Iverson, (2020) in red and Helanow and others, (2021) in black with thickness = 1000 m and slope = 0.25o for $u_{b}$ between 10 m a-1 and 400 m a-1.
Appendix B Appendix B: Reasoning behind line placements in Fig. 2.
@@ Appendix A Appendix A
Form drag, sediment from Zoet and Iverson, (2020) and Minchew and Joughin, (2020) based on the smallest clast sizes used in Zoet and Iverson, (2020) but assumed to continue to sub-milimetre scales (the ice crystal size used is milimetre-scale) and extended to a reasonable but conservative upper limit for subglacial debris size. Skin friction, sediment from Zoet and Iverson, (2020) and Minchew and Joughin, (2020) based on the size of the shear-ring apparatus used in Zoet and Iverson, (2020) and extending to sub milimetre-scale of subglacial till sample used. Assumed to continue to sub-milimetre scales and to be valid at higher spatial scales given low topograhic variation. Weertman sliding from Weertman, (1957) with 1 cm to 10 m scale clearly stated in the text. Budd smaller-scale segment from laboratory tests in Budd and others, (1979) and larger-scale segment from application to west Antarctica in Budd and others, (1984). Nye-Kamb sliding Based on landscape used in Fig. 2 of Nye, (1970). Rate-and-state, Argentière using assumed seismic rupture length in Helmstetter and others, (2015). Cavitation theory from Lliboutry, (1968), Gagliardini and others, (2007), and Helanow and others, (2021) amongst many others. Cavitation theory is often non-dimensionalised but has not exceeded the 25 m scale used in Helanow and others, (2021). The dashed arrow is extended to 100 m to denote a plausible but untested upper limit to cavitation. Intermediate scale processes taken from the lower resolution and upper domain size used in Law and others, (2023) and extended upwards with a dashed arrow to indicate an untested upper limit for the processes described. Argentière Wheel based on the Argentière wheel experiments of Vivian and Bocquet, (1973), Gimbert and others, (2021) and Gilbert and others, (2023) amongst others where the glacier width is $\sim$300 m and 20 m is roughly double the cavity length. Skin drag from Kyrke-Smith and others, (2018) based on the lowermost resolution of radar data (40 m) to the 5 km upper limit of Bedmap2 (Fretwell and others,, 2013). Form drag, topographic from Bingham and others, (2017) and Kyrke-Smith and others, (2018) using a lower radar data resolution of 40-100 m and an upper domain area of 20 km. ’Unified’, from 1 km lower limit suggested in Tsai and others, (2021) to a plausible upper limit of 10 km. Rate-and-state, Siple Coast Taken as broad area over which seismicity is recorded at Siple Coast in Fig. 4 of Podolskiy and Walter, (2016) in contrast to single events of Helmstetter and others, (2015). Form drag, resolution/scale dependent based on arguments within this paper (Section 3.3). Typical glacier and ice sheet model resolution based on standard resolutions from detailed glacier studies to coarse resolution paleo simulations. Typical basal ice crystal size from Thorsteinsson and others, (1995) and Cook and others, (2007).
Appendix C Appendix C
@@
Figure 13: Appendix C. Schematic of a set of sliding processes operating within an inner flow. Sub-processes in cubes should be recognisable from Fig. 4. Red arrows in main panel represent variable basal slip rate and blue arrows represent form drag. Grey mesh with nodes represents model grid cells at the basal boundary.
Appendix D Appendix D

### v3

@@ 6.4 Threshold velocities for regularised-Coulomb sliding
In power-law sliding, two parameters ($C$ and $m$) control the relationship, but even in the simplest regularised-Coulomb relationship at least one further parameter is required ($u_{t}$ in our usage in Table 1) to set the transition point to bounded traction. Laboratory experiments over sediment indicate a transition to bounded traction occurs at around 50 m a-1 (Zoet and Iverson,, 2020) in a controlled setting, while the transition for cavitation over hard beds at small scales (<25 m) is more variable, and possibly as low as 5 m a-1 (Helanow and others,, 2021). Given that uniformly smooth soft beds of consistent material properties, or hard beds of consistent geometry, are somewhat unlikely, it becomes complicated to suggest a reasonable physically based $u_{t}$ value that can be universally applied. For example, Joughin and others, (2019) set a threshold velocity of 300 m a-1 for Pine Island Glacier, Antarctica which sets a clear spatial boundary for the region of bounded traction, but without process-based reasoning for the threshold velocity value. This may be appropriate for the Pine Island setting, but it’s not clear if a 300 m a-1 limit would be equally effective in, for example, rough regions of the Greenland Ice Sheet. As bounded traction significantly increases the sensitivity of outlet glaciers to climate forcing (Ritz and others,, 2015; Tsai and others,, 2015; Brondex and others,, 2019), further work determining the factors controlling the threshold velocity (Trevers and others,, 2024) is necessary. We further note that in marine-terminating settings it may be difficult to distinguish whether the $u_{t}$ value used results in a good match with observations because of an intrinsic switch in processes beyond the threshold velocity, or as a result of the increasing influence of the ocean on subglacial hydrology, with ice velocity acting as a proxy for proximity to the coast.
6.5 Which basal sliding relationship to use?
@@
No single extant sliding relationship accounts for the full breadth of processes or observations that comprise glacier and ice-sheet sliding (Sections 3, 5), making it challenging to recommend a single-tunable-coefficient sliding relationship. Pending future investigations, we make the following suggestions based primarily on heuristic studies (Section 5), whose empirically-validated nature makes them more suitable for application to glacier and ice-sheet models.
We suggest that power-law sliding (or equivalently, pseudo-plastic, Weertman-type sliding), with a setting-dependent $m$, presents a reasonable first choice. Given information from the limited number of studies available (Table 2), and in broad agreement with our discussion (Sections 6.1-6.3), lower $m$ values ($\leq 4$) are then better suited for rougher settings while higher $m$ values ($\geq 5$) are better suited for regions where topography is expected or known to be extensively smooth. Therefore, while $C$ is the main object of the inversion procedure, $m$ may be assigned based on additional constraints, such as topographic roughness. While we do not doubt the efficacy of regularised-Coulomb in some settings (Table 2), it is simpler to manipulate a power-law relationship to approximate regularised-Coulomb using a high value of $m$ than vice versa. As models for glaciers or ice sheets frequently invert for traction coefficients over regions of inferred frozen beds (Goelzer and others,, 2020; Seroussi and others,, 2020), it is further likely that a regularised-Coulomb relationship would misrepresent these areas.
@@
If hydrology is incorporated, the simplified implementation of Budd sliding as used in Choi and others, (2022) and set as the default for ISSM (Larour and others,, 2012), which is functionally the same as power-law sliding but with $N$ included as a multiplication factor is a good starting point. A simplified hydrological model as used in Choi and others, (2022) based on elevation above sea level may then be reasonable over annual timescales (Section 4), though we stress that this is an extremely pressing area for future research – as is the exact way in which $N$ should be incorporated.
@@ 6.5.1 A note on terminology
We hope that this paper provides a clear description of the inner flow sub-processes that a sliding relationship will account for under ideal circumstances, and may also ease some of the angst experienced by modelling practitioners (ourselves included) in labelling bed properties. When we invert for basal properties we are not solely inverting for ‘roughness’, ‘till friction angle’, or ‘friction’ at the ice-bed interface itself – we are likely inverting for all of these processes, and several more, simultaneously. We use ‘ideal circumstances’ here because the traction field is often one of the few tunable parameters in an ice-sheet model and so will in practice also account for so-called ‘compensating errors’ arising from incorrectly prescribed englacial temperature, ice viscosity, subglacial topography, and surface and basal mass balance (Berends and others,, 2023). Last, while we use ‘glacier sliding’ throughout this paper to remain in keeping with its common usage, but make a distinction with ‘slip’ to refer strictly to ice-bed displacement; ‘glacier basal motion’ (e.g. Waller,, 2001; Hubbard,, 2002; Law and others,, 2023) is probably a more precise term.
@@ 7 Conclusions
Glacier sliding is not a straightforward topic. Here, we have done our best to integrate existing understanding into a relatively straightforward framework. This is centred upon an inner flow (Fig. 5) comprised of potentially numerous sliding sub-processes (Fig. 4), which can be represented over a given region through Eq. 8. While it may be possible to reduce this complexity into a relatively simple aggregate equation (such as Eq. 21), we would advocate for the use of a power-law relationship, with careful selection of the $m$ parameter based on the setting, due to its flexibility in representing both regularised-Coulomb and power-law end members.
Our coverage also highlights some clear priorities for future work. Principally, we still do not have a clear methodology for linking stick-slip sliding with continuous sliding relationships, and major questions remain regarding how best to incorporate $N$ within sliding relationships, and how it may be expected to change at the decadal to centennial timescales relevant to projections of ice-sheet and glacier mass loss. The landscape factors influencing the balance of slip, deformation, and geometric resistance also remain unexplored in a quantitative manner, as does the importance and formation, or lack thereof, of subglacial cavitation at large scales. Overall, we hope that the framework presented here allows for a transparent approach to understanding glacier sliding, applying glacier sliding relationships to numerical models, and for directing future research.
Appendix A1 Reasoning behind line placements in Fig. 2.
@@
Form drag, sediment from Zoet and Iverson, (2020) and Minchew and Joughin, (2020b) based on the smallest clast sizes used in Zoet and Iverson, (2020) but assumed to continue to sub-milimetre scales (the ice crystal size used is milimetre-scale) and extended to a reasonable but conservative upper limit for subglacial debris size. Skin friction, sediment from Zoet and Iverson, (2020) and Minchew and Joughin, (2020b) based on the size of the shear-ring apparatus used in Zoet and Iverson, (2020) and extending to sub milimetre-scale of subglacial till sample used. Assumed to continue to sub-milimetre scales and to be valid at higher spatial scales given low topograhic variation. Weertman sliding from Weertman, (1957) with 1 cm to 10 m scale clearly stated in the text. Budd smaller-scale segment from laboratory tests in Budd and others, (1979) and larger-scale segment from application to west Antarctica in Budd and others, (1984). Nye-Kamb sliding Based on landscape used in Fig. 2 of Nye, (1970). Rate-and-state, Argentière using assumed seismic rupture length in Helmstetter and others, (2015). Cavitation theory from Lliboutry, (1968), Gagliardini and others, (2007), and Helanow and others, (2021) amongst many others. Cavitation theory is often non-dimensionalised but has not exceeded the 25 m scale used in Helanow and others, (2021). The dashed arrow is extended to 100 m to denote a plausible but untested upper limit to cavitation. Intermediate scale processes taken from the lower resolution and upper domain size used in Law and others, (2023) and extended upwards with a dashed arrow to indicate an untested upper limit for the processes described. Argentière Wheel based on the Argentière wheel experiments of Vivian and Bocquet, (1973), Gimbert and others, (2021) and Gilbert and others, (2023) amongst others where the glacier width is $\sim$300 m and 20 m is roughly double the cavity length. Skin drag from Kyrke-Smith and others, (2018) based on the lowermost resolution of radar data (40 m) to the 5 km upper limit of Bedmap2 (Fretwell and others,, 2013). Form drag, topographic from Bingham and others, (2017) and Kyrke-Smith and others, (2018) using a lower radar data resolution of 40-100 m and an upper domain area of 20 km. ’Unified’, from 1 km lower limit suggested in Tsai and others, (2021) to a plausible upper limit of 10 km. Rate-and-state, Siple Coast Taken as broad area over which seismicity is recorded at Siple Coast in Fig. 4 of Podolskiy and Walter, (2016) in contrast to single events of Helmstetter and others, (2015). Form drag, resolution/scale dependent based on arguments within this paper (Section 3.3). Typical glacier and ice sheet model resolution based on standard resolutions from detailed glacier studies to coarse resolution paleo simulations. Typical basal ice crystal size from Thorsteinsson and others, (1995) and Cook and others, (2007).
@@
Figure A2: Schematic of a set of sliding processes operating within an inner flow. Sub-processes in cubes should be recognisable from Fig. 4. Red arrows in main panel represent variable basal slip rate and blue arrows represent form drag. Grey mesh with nodes represents model grid cells at the basal boundary.
Appendix A2 Pressure drag from a spherical obstacle
Taking a power law $\tau=M\dot{\epsilon}^{p}$ where $M$ is the consistency index, with an effective viscosity $\mu_{eff}=M\dot{\epsilon}^{p-1}$ and using $\dot{\epsilon}\sim V/L$ where $V$ is the far-field velocity and $L$ is the obstacle size gives
$\mu_{eff}\sim M\frac{V}{L}^{p-1}.$ (23)
Separately, taking the resistance, $F_{D}$, from the normal forces acting on the sphere as
$F_{D}\propto\mu_{eff}VL$ (24)
from Stoke’s law (e.g. Happel and Brenner,, 1983, 4-17.28) and substituting Eq. 23 gives $F_{D}\sim MV^{p}L^{2-p}$ which then gives Eq. 12 if we set $p=1/n$ and $M=K_{R_{i}}$.
Appendix A3 Example expansion of $\boldsymbol{\mathcal{S}_{T}}$
$\boldsymbol{\mathcal{S}}_{T}(x,y)=\begin{cases}\boldsymbol{\mathcal{S}}_{R_{\alpha}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\alpha}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\beta}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\beta}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\gamma}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\gamma}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\alpha}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\beta}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\alpha}\cap R_{\beta}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\alpha}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\gamma}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\alpha}\cap R_{\gamma}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\beta}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\gamma}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\beta}\cap R_{\gamma}\text{ only},\\ \boldsymbol{\mathcal{S}}_{R_{\alpha}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\beta}}(\boldsymbol{u}_{s})+\boldsymbol{\mathcal{S}}_{R_{\gamma}}(\boldsymbol{u}_{s})&\text{if }(x,y)\in R_{\alpha}\cap R_{\beta}\cap R_{\gamma}.\end{cases}$ (25)
Appendix A4 Sliding relationships in geodynamics
@@
Figure A3: A spatial bound to Iken’s bound? a Cavitation (blue) occurring where topography is sufficiently small adapted from Schoof, (2005). b Cavitation can occur at a sufficiently small scale (yellow highlight) but will not occur in larger wavelength depressions. z is height from baseline and x is distance along flow. The thick black line is the glacier bed. $R$ represents a possible region in consideration for the sliding relationship. This figure is drawn in two dimensions for comparison with previous studies, but the same principles apply in three dimensions with realistic topography.